If the area of the t-distribution to the right is 0.01, then find the following:

a. the area to the left of the distribution.

b. the degree of freedom when n=16.

c. the t-value when the confidence level is 99% and df=16


1
Expert's answer
2022-03-28T07:29:35-0400

a. 1-0.01=0.99

b.degree of freedom = n-1=16-1=15

c.2.583